Two vehicles A (mass 520kg) and B (1.7 tons) collide at an intersection. Before the collision, vehicle A traveled at a velocity of 120km/h in an eastern direction and vehicle B traveled with a velocity of 85km/h in a northern direction. The collision is in elastic.
1. Calculate the momentum of Vehicle A before the collision (magnitude only).
2. Calculate the momentum of vehicle B before the collision (magnitude only).
3. Calculate the momentum before the collision (magnitude only).
1
Expert's answer
2020-09-14T10:20:01-0400

By definition, the momentum is:


"p = mv"

where "m" is the mass and "v" is the speed of a vehicle.

1.  The momentum of Vehicle A before the collision


"p_A = m_Av_A = 520kg\cdot 120 km/h = 520kg\cdot 33.(3)m/s \approx 17333 kg\cdot m/s"

2.  The momentum of Vehicle B before the collision


"p_B = m_Bv_B = 1.7\space tons \cdot 85\space km/h = 1700kg\cdot 23.6(1)m/s \approx 40139 kg\cdot m/s"

3.  The momentum before the collision

As far as "p_A" and "p_B" are perpendicular to each other, the magnitude of their sum can be found using the Pythagorean theorem:


"p = \sqrt{p_A^2 + p_B^2} =\sqrt{17333^2 + 40139^2}\approx 43722kg\cdot m/s"

Answer. 1. 17333 kg*m/s, 2. 40139 kg*m/s, 3. 43722 kg*m/s.